# A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?

### v4

@@ IX-A Data-Driven Roadmap of Optimizing LLMs for NL2SQL
Condition 2: Data Volume. For open-source LLMs, optimization is possible in both training and inference phases, while closed-source LLMs allow only inference-stage optimization due to limited access. With extensive nl2sql data, pre-training enhances performance; fine-tuning is suitable for datasets with hundreds to thousands of (nl, sql) pairs. In low-data scenarios, few-shot learning is recommended, while zero-shot methods are essential when labeled data is unavailable. Hardware resources and API costs are also important considerations in selecting the best optimization strategy.
IX-B Decision Flow of Selecting NL2SQL Modules
In Figure 10(b), we present recommendations for choosing nl2sql modules based on specific scenarios, highlighting both benefits and trade-offs. Below, we outline two examples.
@@ IX-B Decision Flow of Selecting NL2SQL Modules
In summary, while each module offers unique advantages for specific nl2sql scenarios, it is essential to balance these benefits with the potential drawbacks in system design.
X Open Problems
@@
Open NL2SQL Problem. In real-world scenarios like government open data platforms, citizens may ask questions that require querying multiple databases and aggregating results. For example, a citizen might ask, “What is the average processing time for tax returns in the last five years?”. Answering this question involves retrieving relevant tables from different databases, such as tax records, processing logs, and statistical reports, and then generating multiple sql queries over them.
@@
The Open nl2sql problem requires translating nl queries into sql queries that access not just a single fixed database but also identify relevant databases (or tables) within a large-scale DBMS. Unlike traditional nl2sql, where a single target database is specified by the user, Open nl2sql may need to generate multiple sql queries that access different databases for a single user question. Thus, the Open nl2sql problem introduces unique challenges, including: (1) database retrieval: accurately identifying and retrieving relevant databases from a vast array of data sources; (2) handling heterogeneous schemas: integrating data with varied structures and terminologies, requiring advanced schema matching and linking techniques; (3) answer aggregation: inferring final answers from multiple sql queries across databases, which demands methods to plan query order, resolve conflicts, and ensure consistency; (4) domain adaptation: generalizing models across domains to address differences in terminology and structure; (5) scalability and efficiency: managing large data volumes while maintaining performance; and (6) evaluating and benchmarking: developing metrics and datasets that accurately reflect real-world complexity for open nl2sql solutions.
@@
Develop Cost-effective NL2SQL Methods. LLM-based nl2sql methods show great potential but are limited by high token consumption, leading to increased costs and slower inference times. In contrast, PLM-based nl2sql methods excel at handling complex sql queries and accurately interpreting database schemas. A promising approach is to combine the strengths of both, developing modular nl2sql solutions or using a multi-agent framework to integrate LLMs and PLMs for the nl2sql task. This hybrid strategy can manage complex queries while conserving tokens and reducing costs. However, effectively balancing LLM and PLM usage to optimize performance and resource efficiency remains an open challenge.
@@
Make NL2SQL Solutions Trustworthy. Ensuring nl2sql solutions are trustworthy is essential for generating accurate and reliable sql, mitigating risk, and reducing the need for manual intervention. Key topics include the following:
@@
Interpreting NL2SQL Solutions. Understanding the reasoning behind a nl2sql model’s performance enhances confidence in its reliability. Explainable AI techniques [153, 154], such as surrogate models [155] and saliency maps [156], can help reveal the decision-making processes within the model. However, their effectiveness in nl2sql contexts, especially with combined LLM and PLM, remains an open question. In addition, multi-agent frameworks using LLMs [157] can improve reliability by dividing the nl2sql task into specialized sub-tasks, each handled by an agent optimized for a particular function. This approach leverages the strengths of multiple agents to enhance overall robustness and accuracy. Nevertheless, effectively coordinating these agents to work harmoniously and optimize their combined performance poses significant challenges in nl2sql contexts.
@@
NL2SQL Debugging Tools. Inspired by code compilers with step-by-step debugging features, a visual debugger for nl2sql could improve accuracy and reliability by measuring semantic and syntactic errors in generated sql queries. Such tools would detect potential errors, enable users to examine the sql generation process, identify mismatches, and understand the logic behind generated sql. However, while traditional compilers capture syntactic errors, nl2sql debugging tools would also need to address semantic errors, i.e., ensuring that generated sql align with the intent of the original nl query.
@@
Interactive NL2SQL Tools. These tools are essential for empowering professional users (e.g., DBAs) to create complex sql queries that span multiple databases, often exceeding 50 lines of code. A key feature is the model’s ability to decompose complex queries into manageable sub-queries, reducing cognitive load and enabling DBAs to focus on each part before reassembling them. Supporting both bottom-up and top-down approaches, these tools allow DBAs to build queries according to their workflow preferences and to correct and refine generated sql as needed. This collaborative process aligns complex sql output with user intent, combining the model’s capabilities with the DBA’s expertise.
@@
Adaptive Training Data Synthesis. Current learning-based nl2sql methods struggle to adapt to new, unseen domains, highlighting a gap in generalizability across data domains. In addition, the performance of these methods heavily relies on the quality, coverage, diversity, and amount of the training data. Therefore, an interesting research problem is to automatically and incrementally generate (nlsql) pairs based on the model performance. The key idea is to dynamically synthesize or augment the training data by leveraging feedback from nl2sql evaluation results. Specifically, by incorporating insights from evaluation metrics and evaluation results, we can identify specific weaknesses of the model. Using this information, we can synthesize training data that continually evolves with the help of LLMs to cover a broader range of scenarios and domains.
@@ XI Conclusion
In this paper, we provide a comprehensive review of nl2sql techniques from a lifecycle perspective in the era of LLMs. We formally state the nl2sql task, discuss its challenges, and present a detailed taxonomy of solutions based on the language models they rely on. We summarize the key modules of language model-powered nl2sql methods, covering pre-processing strategies, translation models, and post-processing techniques. We also analyze nl2sql benchmarks and evaluation metrics, highlighting their characteristics and typical errors. Furthermore, we outline a roadmap for practitioners to adapt LLMs for nl2sql solutions. Finally, we maintain an online nl2sql handbook to guide researchers and practitioners in the latest nl2sql advancements and discuss the research challenges and open problems for nl2sql.

### v5

@@ IX-A Data-Driven Roadmap for Text-to-SQL
Condition 2: Data Volume. For open-source LLMs, optimization is possible in both training and inference phases, while closed-source LLMs allow only inference-stage optimization due to limited access. With extensive Text-to-SQL data, pre-training enhances performance; fine-tuning is suitable for datasets with hundreds to thousands of (nl, sql) pairs. In low-data scenarios, few-shot learning is recommended, while zero-shot methods are essential when labeled data is unavailable. Hardware resources and API costs are also important considerations in selecting the best optimization strategy.
IX-B Decision Flow of Selecting Text-to-SQL Modules
In Figure 10(b), we present recommendations for choosing Text-to-SQL modules based on specific scenarios, highlighting both benefits and trade-offs. Below, we outline two examples.
@@ IX-B Decision Flow of Selecting Text-to-SQL Modules
In summary, while each module offers unique advantages for specific Text-to-SQL scenarios, it is essential to balance these benefits with the potential drawbacks in system design.
X Limitations and Open Problems
We analyze the limitations of LLM-based methods and propose corresponding open problems, highlighting unresolved challenges and suggesting directions for future research.
Limitations of Current LLM-based Solutions. Although recent LLM-based Text-to-SQL methods have made significant progress, they still face several challenges when dealing with complex queries in real-world scenarios. First, existing methods are typically trained and executed on a single, fixed database, which limits their ability to handle open environments that require cross-database queries and multi-source data aggregation. Second, although LLMs have strong natural language understanding capabilities, they incur high token consumption during inference, leading to high costs and low efficiency. In addition, most Text-to-SQL methods lack interpretability and debugging mechanisms, making it difficult for users to understand how the model generates SQL or to detect and fix potential semantic errors. Finally, current methods show limited adaptability to new domains and rely heavily on high-quality training data; how to automatically generate targeted training samples based on model feedback remains an open problem. These limitations reveal shortcomings in cross-database scalability, inference efficiency, system reliability, and data adaptability, highlighting the need for more efficient, trustworthy, and scalable Text-to-SQL solutions.
@@ X Limitations and Open Problems
Open-Domain Text-to-SQL Problem. In real-world scenarios like government open data platforms, citizens may ask questions that require querying multiple databases and aggregating results. For example, answering “What is the average processing time for tax returns in the last five years?” requires retrieving tables from multiple databases (e.g., tax records, processing logs, and statistical reports) and generating multiple sql queries over them. Unlike traditional Text-to-SQL, where a single target database is specified by the user, Open Text-to-SQL may need to generate multiple sql queries that access different databases for a single nl.
@@
Thus, the Open Text-to-SQL problem introduces unique challenges, including: (1) database retrieval: accurately identifying and retrieving relevant databases from a vast array of data sources; (2) handling heterogeneous schemas: integrating data with varied structures and terminologies, requiring advanced schema matching and linking techniques; (3) answer aggregation: inferring final answers from multiple sql queries across databases, which demands methods to plan query order, resolve conflicts, and ensure consistency; (4) domain adaptation: generalizing models across domains to address differences in terminology and structure; (5) scalability and efficiency: managing large data volumes while maintaining performance; and (6) evaluating and benchmarking: developing metrics and datasets that accurately reflect real-world complexity for Open Text-to-SQL solutions.
@@
Develop Cost-effective Text-to-SQL Methods. LLM-based Text-to-SQL methods show great potential but are limited by high token consumption, leading to increased costs and slower inference times. In contrast, PLM-based Text-to-SQL methods excel at handling complex sql queries and accurately interpreting database schemas. A promising approach is to combine the strengths of both, developing modular Text-to-SQL solutions or using a multi-agent framework to integrate LLMs and PLMs for the Text-to-SQL task (as shown in Table III). In parallel, efforts have aimed to improve LLM-based efficiency. EllieSQL [155] employs complexity-aware routing to enhance cost-efficiency by assigning queries to suitable LLM-based generators.
TABLE III: Resource Consumption Statistics on the Spider.
Method Base Model API Tokens/SQL Latency/SQL (s)
RESDSQL [7] PLM - 1.91
RESDSQL+NatSQL [7] PLM - 1.97
ZeroNet [44] PLM+LLM 377 3.72
DIN-SQL [5] LLM 3579 10.34
@@
Make Text-to-SQL Solutions Trustworthy. Ensuring Text-to-SQL solutions are trustworthy is essential for generating accurate and reliable sql, mitigating risk, and reducing the need for manual intervention. Topics include the following:
@@
Interpreting Text-to-SQL Solutions. Understanding the reasoning behind a Text-to-SQL model’s performance enhances confidence in its reliability. Explainable AI techniques [156, 157], such as surrogate models [158] and saliency maps [159], aim to reveal model decisions. However, their effectiveness in Text-to-SQL contexts, especially with combined LLM and PLM, remains an open question. In addition, multi-agent LLM frameworks [160] improve reliability by splitting Text-to-SQL into specialized sub-tasks. Although this approach improves robustness, coordinating agents to ensure consistent and optimized performance remains a major challenge.
@@
Text-to-SQL Debugging Tools. Inspired by compiler design, a debugger for Text-to-SQL could improve accuracy and reliability by measuring semantic and syntactic errors in generated sql queries. Such tools would detect potential errors, enable users to examine the sql generation process and identify mismatches [33, 34]. However, achieving this goal presents significant challenges. Traditional code compilers primarily capture syntactic errors, while Text-to-SQL debugging must also address semantic errors, i.e., ensuring that the generated sql query accurately reflects the intent of the nl query.
@@
Interactive Text-to-SQL Tools. These tools are essential for empowering professional users (e.g., DBAs) to create complex sql queries that span multiple databases, often exceeding 50 lines of code. A key feature is the model’s ability to decompose complex queries into manageable sub-queries, reducing cognitive load and enabling DBAs to focus on each part before reassembling them. Supporting both bottom-up and top-down workflows, such tools enable users to iteratively refine outputs, align sql generation with intent, and integrate model assistance with domain expertise.
@@
Adaptive Training Data Synthesis. Learning-based Text-to-SQL models often fail to generalize to unseen domains, partly due to limited training data coverage, quality, and diversity. Therefore, an interesting research problem is to automatically and incrementally generate (nlsql) pairs based on the model performance. Specifically, by incorporating insights from evaluation metrics and evaluation results, we can identify specific weaknesses of the model. Using this information, we can synthesize training data that continually evolves with the help of LLMs to cover a broader range of domains.
@@ XI Conclusion
In this paper, we comprehensively review Text-to-SQL techniques from a lifecycle perspective in the LLM era. We formally define the Text-to-SQL task, discuss key challenges, and propose a taxonomy based on underlying language models. We summarize key modules of language model-driven methods, including pre-processing, translation, and post-processing strategies. Furthermore, we analyze benchmarks and evaluation metrics, highlighting their characteristics and common errors. We also offer a practical roadmap for adapting LLMs to Text-to-SQL tasks and maintain an online handbook with the latest advancements, discussing ongoing challenges and open problems.